What mass of carbon dioxide can be expected from a complete combustion of 1000.00 kg of propane with 1000.00 kg of oxygen?
If the percent yield for your reaction and collection technique is 72.5%, then how much carbon dioxide was actually created?
1
Expert's answer
2020-06-18T14:44:57-0400

combustion reaction for propane:

C3H8+5O2--->3CO2+4H2O


moles of propane taken = 1000000g/44gmol-1

= 22727.27 moles

moles of oxygen taken = 1000000/32

= 31250 moles


here limiting reagent is oxygen so whole oxygen will be consumed.


no. of moles of carbon dioxide formed = (3/5)*31250

=18750 moles


since yield is 75% so obtained CO2 is (75/100)*18750

= 14062.5 moles


mass of CO2 obtained= 14062.5*44 g

=618750 grams = 618.750 kg